# From molecules to scaffolds to functional groups: building context-dependent molecular representation via multi-channel learning

### v2

@@ 4 Methods
The ROGI value has been shown to be strongly correlated with other metrics (e.g., SARI), as well as the representation performance on the given dataset [2]. One of the main advantages of using ROGI is that it is a generalized QSPR metric applicable to a wide range of molecular representations, including both fingerprints and neural representations, while other metrics like SARI are tailored only for molecular fingerprints. This allows us to compute landscape roughness on pre-trained molecular representations and perform deeper QSPR analysis.
@@
Instead of using the same readout operations as the conventional GNNs, we adopt the prompt-guided aggregation, which is achieved using the multi-head attention. As shown in Figure 1e, the embedding of the prompt token is treated as the attention query, while the representations of nodes/atoms , are viewed as the keys and values. The resulted prompt-aware graph representation , where the attention weight for each node is computed as , and , , and . Here represent parameters in the linear transformations, and is the scaling factor. Essentially, the prompt-aware is aggregated from the weighted average of .
@@
where denotes the latent representation for sample , and being the representation for its augmentation. The function measures the L2 distance between two vectors. In general, this objective enforces the pair-wise distancing difference between <, > and <, > to be at least margin. However, as we discuss before, this formulation can only lead to a coarse-grained representation space, while neglecting the existing structural relationship among molecules (e.g., shared rings or functional groups). Also, this formulation does not pose any constraints on the representation space beyond the margin (see Supplementary Information for example illustration). Therefore, we propose to add an additional contrastive formulation with adaptive margin among negative samples and the anchor.
@@
At the fine-tune stage, the pre-trained model parameter are used to initialize the model. Besides, we introduce an additional prompt selection module to combine representations from different channels into the task-specific (i.e., context-dependent) composite representation. Essentially, it learns the relevance between different pre-trained molecular knowledge and the downstream application, hence bridging the gap between pre-training and fine-tuning objectives. To incorporate task-specific SPR information into the model at an early stage of fine-tuning, we propose to initialize the prompt weights from computing the roughness index (ROGI) [2]. In short, the initial prompt weights should lead to a composite representation with the lowest ROGI value (i.e., smoothest quantitative structure-property landscape) with respect to the current application. A low ROGI value indicates smoother landscape, hence better modellability. We use a simple Bayesian optimization pipeline to find the best initialization with the lowest ROGI value on the training set. Essentially, the input parameters of the Bayesian optimization are learnable scalers, where is the number of channels. The black-box utility function first computes the distribution over values by treating the input scalars as logits, and calculates the composite representation as well as its corresponding roughness index. We utilize the Quasi MC-based batch Expected Improvement as the acquisition function. After the initialization, the entire model except for the prompt-guided aggregation module is optimized towards the molecular property labels. The parameters in the aggregation modules are fixed during fine-tuning as we aim to use it as a pooling layer independent of the downstream applications. Hypothetically, the node attention scores within the aggregation module should reflect the expected atom importance (e.g., SCD’s aggregation attention captures only the scaffold atoms). See Supplementary Information for more discussion and case studies in prompt-guided node aggregations.
@@
We pre-train our framework using the molecules from ZINC15 [42], which is an open-sourced database that contains 2 million unlabelled drug-like compounds. We use CReM [34], an open-sourced molecule mutation framework, for the scaffold-invariant molecule perturbation. To be more specific, we first use RDKit [1] to identify the Bemis-Murcko scaffold of molecules, as well as the connection sites (i.e., atom indices) between the scaffold and the side chains. The CReM algorithm then takes these indices and performs fragment replacement using an external fragment pool. Perturbation results are further filtered by chemical validity and the maximum number of changed atoms allowed. At last, we are able to successfully compute perturbations for 1,882,537 out of 2 million molecules. These molecules form our final pre-train database. In the contrastive-based pre-training, we consider five positive samples for each anchor molecule. In molecule distancing, we randomly apply subgraph masking five times on the original molecule. In scaffold distancing, we randomly sample five scaffold-invariant perturbations. In terms of the model architecture, we follow the same architecture setup as the graph neural network in [60, 22, 52, 25], which is a 5-layer Graph Isomorphism Network (GIN) [59] with hidden dimension size equals to 300. In addition, we also pre-train a GPS Graph Transformer model [36] with the same number of layer and hidden dimension size. We use the same molecule feature sets as works in [38, 14]. During fine-tuning, we utilize the scaffold split method [23] for MoleculeNet [56], with a train, validation, and test ratio of 8:1:1. On the other hand, we apply the stratified split for MoleculeACE [50], as proposed by itself, with a train, validation, and test ratio of 8:1:1.
@@
We consider eleven baseline pre-training methods in our experiments. 1. GraphLoG [60] achieves a hierarchical prototypical embedding space with the conventional graph perturbation techniques; 2. D-SLA [25] proposes the discrepancy learning to refine the embedding space with the conventional graph perturbation techniques; 3. MolCLR [52] utilizes the NT-Xent [6] contrastive loss with molecule perturbation techniques, including atom/bond editing and subgraph masking; 4. KANO [14] introduces the knowledge graph prompting techniques to augment molecular graph, while the augmentations are also learned in the contrastive manner. 5. MoLFormer [39] performs large scale of masked language modeling on SMILES sequence using a linear attention Transformer. 6. Hu et. al [22] adopts masked attribute prediction and molecular subgraph prediction; 7. KPGT [27] proposes the Line Graph Transformer on molecular graphs and performs knowledge prediction. It takes a masked molecular graph, molecule fingerprint, and molecule descriptors as input, and aims to reconstruct the molecule fingerprint and descriptors during pre-training. 8. GROVER [38] proposes the GTransformer architecture and applies context prediction learning on molecular graphs; 9. GEM [13] encodes the 3d geometric information of molecules and predicts the bond angle and atom distance. 10. Uni-Mol [65] also incorporates 3d information of the molecule and pre-trains a SE(3) Transformer [15] using 3d position recovery and masked atom prediction. We further include 11. GraphMVP [28], which performs both contrastive and predictive learning between 2d toplogical and 3d geometric information of the molecule. We also train the GIN [59] and GPS [36] models from scratch for comparison. We reproduce the performance of MolCLR, GROVER, GEM, KPGT, MoLFormer, and KANO on the MoleculeNet benchmark using their provided checkpoints and code repositories. Note that the reported results of GROVER, KPGT, MoLFormer and KANO in their original papers are evaluated under the balanced scaffold split, which is different than the deterministic scaffold split used in this work and the rest of the baseline methods. For the MoleculeACE benchmark, we fine-tune all the baseline models using their corresponding repositories and provided model checkpoints.
@@
We propose to analyze the dynamics of representation space during fine-tuning to evaluate the representation robustness and generalizability. To be more specific, we probe the mapping of the representation space at five different training timestamps (epoch 0, epoch 10, epoch 20, epoch 50, epoch 100). The 2-D view is constructed via the T-SNE [49] dimension reduction technique. Besides the visualization, we also report four additional metrics that capture the representation characteristics. We report the ROGI value and Rand index in the training set, plus the R-squared value and cliff-noncliff distance ratio in the validation set. The Rand index is calculated from two clustering assignments. The first clustering is formed using k-means clustering with the molecule ECFP4 fingerprint (radius=2, nBits=512). The second clustering is formed with using k-means clustering with the representation at the current timestamp. We set the same number of clusters for these two clusterings. In terms of the cliff-noncliff distance ratio, we first identify the cliff and the noncliff molecule pairs using the same principle in MoleculeACE [50]. Then, we compute and average the representation distance of each pair. At last, the cliff-noncliff ratio is formed by normalizing their average distance.

### v3

@@ Methods
The ROGI value has been shown to be strongly correlated with other metrics (e.g., SARI), as well as the representation performance on the given dataset [41]. One of the main advantages of using ROGI is that it is a generalized QSPR metric applicable to a wide range of molecular representations, including both fingerprints and neural representations, while other metrics like SARI are tailored only for molecular fingerprints. This allows us to compute landscape roughness on pre-trained molecular representations (Supplementary Fig. 4 and Supplementary Note 3) and perform deeper QSPR analysis.
@@
Instead of using the same readout operations as the conventional GNNs, we adopt the prompt-guided aggregation, which is achieved using the multi-head attention. As shown in Figure 1e, the embedding of the prompt token is treated as the attention query, while the representations of nodes/atoms , are viewed as the keys and values. The resulted prompt-aware graph representation , where the attention weight for each node is computed as , and , , and . Here represent parameters in the linear transformations, and is the scaling factor. Essentially, the prompt-aware is aggregated from the weighted average of linear projection of .
@@
where denotes the latent representation for sample , and being the representation for its augmentation. The function measures the L2 distance between two vectors. In general, this objective enforces the pair-wise distancing difference between <, > and <, > to be at least margin. However, as we discuss before, this formulation can only lead to a coarse-grained representation space, while neglecting the existing structural relationship among molecules (e.g., shared rings or functional groups). Also, this formulation does not pose any constraints on the representation space beyond the margin (Supplementary Fig. 3). Therefore, we propose to add an additional contrastive formulation with adaptive margin among negative samples and the anchor.
@@
At the finetune stage, the pre-trained model parameter are used to initialize the model. Besides, we introduce an additional prompt selection module to combine representations from different channels into the task-specific (i.e., context-dependent) composite representation. Essentially, it learns the relevance between different pre-trained molecular knowledge and the downstream application, hence bridging the gap between pre-training and fine-tuning objectives. To incorporate task-specific SPR information into the model at an early stage of fine-tuning, we propose to initialize the prompt weights from computing the roughness index (ROGI) [41]. In short, the initial prompt weights should lead to a composite representation with the lowest ROGI value (i.e., smoothest quantitative structure-property landscape) with respect to the current application. A low ROGI value indicates smoother landscape, hence better modellability. We use a simple Bayesian optimization pipeline to find the best initialization with the lowest ROGI value on the training set. Essentially, the input parameters of the Bayesian optimization are learnable scalers, where is the number of channels. The black-box utility function first computes the distribution over values by treating the input scalars as logits, and calculates the composite representation as well as its corresponding roughness index. We utilize the Quasi MC-based batch Expected Improvement as the acquisition function. After the initialization, the entire model except for the prompt-guided aggregation module is optimized towards the molecular property labels. The parameters in the aggregation modules are fixed during fine-tuning as we aim to use it as a pooling layer independent of the downstream applications. As shown in Supplementary Figs. 7-9, the node attention scores within the aggregation module reflect the expected atom importance. For example, the SCD’s aggregation attention captures mostly the scaffold atoms. We further discover that the learned prompt weights exhibit patterns that are well-aligned with the necessary chemical knowledge for solving the downstream tasks (Supplementary Fig. 12, Supplementary Table 1, and Supplementary Note 8).
@@
We pre-train our framework using the molecules from ZINC15 [38], which is an open-sourced database that contains 2 million unlabelled drug-like compounds. We use CReM [64], an open-sourced molecule mutation framework, for the scaffold-invariant molecule perturbation. To be more specific, we first use RDKit [65] to identify the Bemis-Murcko scaffold of molecules, as well as the connection sites (i.e., atom indices) between the scaffold and the side chains. The CReM algorithm then takes these indices and performs fragment replacement using an external fragment pool. Perturbation results are further filtered by chemical validity and the maximum number of changed atoms allowed. At last, we are able to successfully compute perturbations for 1,882,537 out of 2 million molecules. These molecules form our final pre-train database. In the contrastive-based pre-training, we consider five positive samples for each anchor molecule. In molecule distancing, we randomly apply subgraph masking five times on the original molecule. In scaffold distancing, we randomly sample five scaffold-invariant perturbations. In terms of the model architecture, we follow the same architecture setup as the graph neural network in [22, 18, 21, 23], which is a 5-layer Graph Isomorphism Network (GIN) [43] with hidden dimension size equals to 300. In addition, we also pre-train a GPS Graph Transformer model [44] with the same number of layer and hidden dimension size. We use the same molecule feature sets as works in [19, 24]. During fine-tuning, we utilize the scaffold splits method [42] for MoleculeNet [5], with a train, validation, and test ratio of 8:1:1. On the other hand, we apply the stratified splits for MoleculeACE [6], as proposed by itself, with a train, validation, and test ratio of 8:1:1.
@@
We consider twelve baseline pre-training methods in our experiments. 1. GraphLoG [22] achieves a hierarchical prototypical embedding space with the conventional graph perturbation techniques; 2. D-SLA [23] proposes the discrepancy learning to refine the embedding space with the conventional graph perturbation techniques; 3. MolCLR [21] utilizes the NT-Xent [33] contrastive loss with molecule perturbation techniques, including atom/bond editing and subgraph masking; 4. KANO [24] introduces the knowledge graph prompting techniques to augment molecular graph, while the augmentations are also learned in the contrastive manner. 5. MoLFormer [28] performs large scale of masked language modeling on SMILES sequence using a linear attention Transformer. 6. Hu et. al [18] adopts masked attribute prediction and molecular subgraph prediction; 7. KPGT [45] proposes the Line Graph Transformer on molecular graphs and performs knowledge prediction. It takes a masked molecular graph, molecule fingerprint, and molecule descriptors as input, and aims to reconstruct the molecule fingerprint and descriptors during pre-training. 8. GROVER [19] proposes the GTransformer architecture and applies context prediction learning on molecular graphs; 9. GEM [25] encodes the 3D geometric information of molecules and predicts the bond angle and atom distance. 10. Uni-Mol [27] also incorporates 3D information of the molecule and pre-trains a SE(3) Transformer [66] using 3D position recovery and masked atom prediction. We further include 11. GraphMVP [46], which performs both contrastive and predictive learning between 2D toplogical and 3D geometric information of the molecule. 12. ImageMol [47], which leverages visual information of molecule images for molecular property prediction. We also train the GIN [43] and GPS [44] models from scratch for comparison. We reproduce the performance of MolCLR, GROVER, GEM, KPGT, MoLFormer, and KANO on the MoleculeNet benchmark using their provided checkpoints and code repositories. Note that the reported results of GROVER, KPGT, MoLFormer and KANO in their original papers are evaluated under the balanced scaffold splits, which is different than the deterministic scaffold splits used in this work and the rest of the baseline methods. Both splits aim to hold out a set of molecules with scaffolds that are not seen during training. The deterministic scaffold splits prioritizes the satisfaction of the specified split ratio, while the balanced scaffold splits prioritizes the balance of scaffold frequency across the training, validation, and test sets. We use the deterministic scaffold splits in our experiments primarily because most of the baselines also employ it. All other aspects, including hyperparameter choices for both the models and the training setup, remain consistent with the default configurations of these baseline methods for replicating the results. For the MoleculeACE benchmark, we leverage their corresponding repositories, default configurations, and the provided model checkpoints of these baseline methods and fine-tune each task using either MoleculeACE’s stratified splits or random splits.
@@
We propose to analyze the dynamics of representation space during fine-tuning to evaluate the representation robustness and generalizability. To be more specific, we probe the mapping of the representation space at five different training timestamps (epoch 0, epoch 10, epoch 20, epoch 50, epoch 100). The 2-D view is constructed via the T-SNE [67] dimension reduction technique. Besides the visualization, we also report four additional metrics that capture the representation characteristics. We report the ROGI value and Rand index in the training set, plus the R-squared value and cliff-noncliff distance ratio in the validation set. The Rand index is calculated from two clustering assignments. The first clustering is formed using k-means clustering with the molecule ECFP4 fingerprint (radius=2, nBits=512). The second clustering is formed using k-means clustering with the representation at the current timestamp. We set the same number of clusters for these two clusterings. In terms of the cliff-noncliff distance ratio, we first identify the cliff and the noncliff molecule pairs using the same principle in MoleculeACE [6]. Then, we compute and average the representation distance of each pair. At last, the cliff-noncliff ratio is formed by normalizing their average distance.
@@
Data Availability
All relevant data supporting the key findings of this study are publicly available. All datasets utilized in this paper can be found at https://github.com/yuewan2/MolMCL or the Zenodo repository at https://doi.org/10.5281/zenodo.14010436 [68]. This includes the processed pre-train dataset of ZINC15, the MoleculeNet benchmark (originally from https://github.com/deepchem/deepchem) and the MoleculeACE benchmark (originally from https://github.com/molML/MoleculeACE). The reference protein structure for 3L1S used in this study is available in the Protein Data Bank under accession code 3L1S [https://www.rcsb.org/structure/3l1s]. Source data are provided with this paper.
Code Availability
The source code of this work can be accessed via the GitHub repository at https://github.com/yuewan2/MolMCL or the Zenodo repository at https://doi.org/10.5281/zenodo.14010436 [68].
Acknowledgements
We have no specific acknowledgments to report for this work.
Author contributions
Y.W. proposed and implemented the methodology, and conducted the computational experiments. J.W. carried out the chemical analysis of the model’s predictions. T.H., C.H., and X.J. provided guidance throughout the project. Y.W. and J.W. led the paper writing and all authors contributed.
Computing Interests
All authors declare no competing interests.